### v1

# Bias Amplification: Language Models as Increasingly Biased Media

Abstract
As Large Language Models (LLMs) become increasingly integrated into various facets of society, a significant portion of online text consequently become synthetic. This raises concerns about bias amplification, a phenomenon where models trained on synthetic data amplify the pre-existing biases over successive training iterations. Previous literature seldom discusses bias amplification as an independent issue from model collapse. In this work, we address the gap in understanding the bias amplification of LLMs with four main contributions. Firstly, we propose a theoretical framework, defining the necessary and sufficient conditions for its occurrence, and emphasizing that it occurs independently of model collapse. Using statistical simulations with weighted maximum likelihood estimation, we demonstrate the framework and show how bias amplification arises without the sampling and functional form issues that typically drive model collapse. Secondly, we conduct experiments with GPT-2 to empirically demonstrate bias amplification, specifically examining open-ended generational political bias with a benchmark we developed. We observe that GPT-2 exhibits a right-leaning bias in sentence continuation tasks and that the bias progressively increases with iterative fine-tuning on synthetic data generated by previous iterations. Thirdly, we explore three potential mitigation strategies: Overfitting, Preservation, and Accumulation. We find that both Preservation and Accumulation effectively mitigate bias amplification and model collapse. Finally, using novel mechanistic interpretation techniques, we demonstrate that in the GPT-2 experiments, bias amplification and model collapse are driven by distinct sets of neurons, which aligns with our theoretical framework.
Bias Amplification: Language Models as Increasingly Biased Media
Ze Wang1,2, Zekun Wu1,2, Jeremy Zhang3, Navya Jain 2 Xin Guan 1, Adriano Koshiyama 1 1Holistic AI, 2University College London, 3Emory University, Correspondence: ze.wang.19@ucl.ac.uk, zekun.wu@holisticai.com
1 Introduction
Large language models (LLMs) are trained on vast amounts of text scraped from the internet, which plays a crucial role in improving their capabilities, whether through emergent abilities (Wei et al., 2022) or scaling laws (Kaplan et al., 2020). However, as LLMs become more widely integrated into human society—for example, in content creation and summarization in media, academia, and business (Maslej et al., 2024)—concerns are mounting that a significant portion of online text in the future may be generated, either entirely or partially, by LLMs (Peña-Fernández et al., 2023; Porlezza and Ferri, 2022; Nishal and Diakopoulos, 2024). This highlights a significant and underexplored risk: bias amplification, referring to the degradation of fairness within LLMs over self-consuming training loops (Mehrabi et al., 2022; Taori and Hashimoto, 2022), where models progressively amplify the pre-existing biases. This concern initially arises from the tendency of LLMs to learn from biased datasets. For example, Parrish et al. (2022); Wang et al. (2024); Bender et al. (2021) show that LLMs absorb inherent stereotypes, such as racial and gender discrimination, embedded in human-generated text. Additionally, Haller et al. (2023); Rettenberger et al. (2024a) demonstrated that LLMs can be aligned with specific political ideologies by fine-tuning them on biased datasets. Moreover, Wyllie et al. (2024) shows that classifiers trained on synthetic data increasingly favor a particular class label over successive generations, while the shrinking diversity observed by Alemohammad et al. (2023); Hamilton (2024) suggests a risk that certain demographic groups may become progressively underrepresented in the outputs of LLMs.
The amplification of biases has profound societal implications. It can lead to the perpetuation of stereotypes, reinforcement of social inequalities, and the marginalization of underrepresented groups. In the context of political bias, this can influence public opinion, skew democratic processes, and exacerbate polarization. Understanding and mitigating bias amplification is therefore crucial to ensure that LLMs contribute positively to society and do not inadvertently cause harm. Nevertheless, despite the literature on bias amplification in discriminative models, there is a notable lack of comprehensive frameworks and empirical studies specifically addressing bias amplification for LLMs, as shown in Section 2. In contrast to discriminative models, where bias amplification can be attributed to overfitting the dominant features in the training dataset, its occurrence in LLMs may stem from a more nuanced issue. As we will show in Sections 3 and 5, even when the training data is unbiased, the model can still amplify its pre-existing biases during the training process.
In this paper, we seek to fill this research gap by proposing a theoretical framework that establishes and explains the causes of bias amplification in LLMs. Additionally, we perform both statistical simulation and LLM experiments to demonstrate and exemplify the theorem. In summary, our main contributions are as follows:
1. Theoretical Framework: We establish the necessary and sufficient conditions for bias amplification (i.e. Theorem 1). The theorem aids in understanding the cause of bias amplification and in distinguishing between bias amplification and model collapse. We conduct statistical simulations using weighted maximum likelihood estimation to illustrate the theorem (see Section 3.2).
2. Benchmarking Tool: We trained a highly accurate classifier capable of detecting political leaning in long-text content. With this classifier, we offer a benchmark for evaluating political bias in LLMs through open-generation tasks, filling a gap not covered in current bias studies (see Sections 2 and 4.3).
3. Empirical Demonstration and Mitigation Strategies: We demonstrate bias amplification in terms of political bias in GPT-2 using our benchmarking tool: the model exhibits a right-leaning bias in sentence continuation tasks and becomes increasingly right-leaning over successive generations (see Section 5.1). Additionally, we conducted experiments with three potential mitigation strategies, i.e. Overfitting, Preservation, and Accumulation, comparing their effectiveness, and found that some are surprisingly effective (see Section 5.3).
4. Mechanistic Interpretation: Building on our framework, we propose an innovative mechanistic interpretation pipeline that identifies two distinct sets of neurons responsible for bias amplification and model collapse during iterative fine-tuning experiments with GPT-2. We found minimal overlap between these two sets, supporting our theorem that bias amplification can occur independently of model collapse (see Section 5.4).
The rest of the paper is organized as follows: We first discuss related work in Section 2. In Section 3, we present our theoretical framework for bias amplification and conduct a statistical simulation. Section 4 describes the experimental setup for our LLM experiments, including data preparation and model fine-tuning. In Section 5, we present our empirical findings on bias amplification and the effectiveness of mitigation strategies. Finally, in Section 6, we discuss the implications of our work, concluding with limitations in Section 7.
2 Background and Related Work
Bias Amplification. Bias amplification in machine learning models has been seriously studied in the context of visual recognition tasks. For example, Zhao et al. (2017) found that Conditional Random Fields can exacerbate social biases present in the training data. This study proposed an in-process mitigation approach, employing Lagrangian Relaxation to enforce constraints that ensure the model’s bias performance remains closely aligned with the biases in the training data. Following this, Mehrabi et al. (2022) proposed the concept of bias amplification in feedback loops, where biased models not only amplify the bias present in their training data but also interact with the world in ways that generate more biased data for future models. Xu et al. (2023); Zhou et al. (2024) examined bias amplification in recommendation models, showing how these models reinforce their understanding of mainstream user preferences from training data, leading to an overrepresentation of such preferences in historical data and neglecting rarely exposed items—similar to the concept of sampling error discussed in (Shumailov et al., 2024). Additionally, Wyllie et al. (2024); Taori and Hashimoto (2022) demonstrated that classifiers trained on synthetic data increasingly favor specific class labels over successive generations. Likewise, Ferbach et al. (2024); Chen et al. (2024) observed bias amplification in generative models such as Stable Diffusion, characterized by the overrepresentation of features from the training dataset. Moreover, we conducted a comprehensive review of the literature on model collapse which we put in Appendix L.
Political Biases. LLMs become a new form of "media" that people rely on for global news (Maslej et al., 2024), attracting significant research attention to political biases. Rettenberger et al. (2024b); Shumailov et al. (2024); Feng et al. (2024) explored the bias through voting simulations within the spectrum of German political parties, consistently finding a left-leaning bias in models like GPT-3 and Llama3-70B. Similarly, for the U.S. political landscape, Rotaru et al. (2024); Motoki et al. (2024) identified a noticeable left-leaning bias in ChatGPT and Gemini when tasked with rating news content, evaluating sources, or responding to political questionnaires.
3 Theoretical Framework
In this section, we formalize the conditions under which bias amplification occurs during recursive training on synthetic data. This framework is designed to intuitively illustrate the primary causes of bias amplification. Next, we demonstrate bias amplification and the theorem using Weighted Maximum Likelihood Estimation (WMLE).
3.1 The Conditions for Bias Amplification
Bias amplification can arise from two key factors. The first is referred to as bias projection. It arises when the bias projection coefficient (explained below) is negative, indicating that the gradient update during training will fully align with the existing bias in the model parameters, thus amplifying it. Consider a fine-tuning process where the pre-trained model parameters $\theta_{t}$ can be decomposed into unbiased and biased components:
$\theta_{t}=\theta_{t,\text{unbiased}}+\theta_{t,\text{biased}}.$
During gradient-based optimization, the update rule is:
$\theta_{t+1}=\theta_{t}-\eta\nabla_{\theta}\mathcal{L}_{\text{ft}}(\theta_{t}),$
where $\eta$ is the learning rate, and $\mathcal{L}_{\text{ft}}$ denotes the fine-tuning loss function. Substituting the decomposition of $\theta_{t}$ and taking the projection, we have:
$\theta_{t+1}=\theta_{t,\text{unbiased}}+\theta_{t,\text{biased}}-\eta\left(\frac{\mathbf{\theta_{t,\text{biased}}}}{\|\mathbf{\theta_{t,\text{biased}}}\|}\right)c_{t}$
where $c_{t}$ is the bias projection coefficient, measuring the projection of the gradient onto the normalized biased component of the parameters:
$c_{t}=\left(\frac{\theta_{t,\text{biased}}}{\|\theta_{t,\text{biased}}\|}\right)^{\top}\nabla_{\theta}\mathcal{L}_{\text{ft}}(\theta_{t}).$ (1)
If $c_{t}<0$, the gradient update will reinforce the biased component, leading to bias amplification, i.e. $\Delta|\theta_{\text{biased}}|>0$. This occurs because the gradient descent step moves the parameters further in the direction of the existing bias. Intuitively, the bias in the output is influenced by specific neurons, and increasing the weights of these neurons best aligns with the training objective for the given dataset. This does not require the dataset itself to be biased, but rather that it contains features, aside from bias, that align more closely with the functionalities of biased neurons relative to the unbiased neurons considering the loss function.
The second factor is referred to as constraint deficiency. It occurs when the loss or objective function lacks explicit constraints or regularization to counteract bias projection. As a result, the optimization process does not prevent the reinforcement of bias when bias projection is present. To prevent bias amplification theoretically, the loss function can incorporate de-biasing constraints. Consider a modified fine-tuning objective: $\mathcal{L}_{\text{new}}(\theta)=\mathcal{L}_{\text{original}}(\theta)-\lambda\sum_{i=1}^{n}\max(0,c^{i})^{2}+\gamma\sum_{i=1}^{n}\min(0,c^{i})^{2}$, where $c^{i}$ represents the bias projection coefficient for a specific dimension. $\lambda,\gamma\geq 0$ are the Lagrangian multipliers for the two constraints:
(1) $\sum_{i=1}^{n}\max(0,c^{i})^{2}\geq 0$,   i.e. encouraging positive components.
(2) $\sum_{i=1}^{n}\min(0,c^{i})^{2}\leq 0$,  i.e. eliminating negative components.
The absence of such constraints allows bias amplification when bias projection is present. Based on these findings, we establish the following:
Theorem 1 (Condition for Bias Amplification).
Bias amplification, a self-reinforcing process where a model trained on synthetic data amplifies the pre-existing biases from previous training, leading to an increased bias in the data it generates for future training, occurs if both bias projection and constraint deficiency are present. It occurs only if there is bias projection in at least one parameter dimension and constraint deficiency is present.
Alongside understanding the causes of bias amplification, this also implies that bias amplification can occur even without model collapse. Specifically, statistical approximation and functional expressivity errors [^1] We have not discuss functional approximation error here, as it is a general concept that encompasses any structural setting, from learning to objective function, that may lead to suboptimal training outcomes, making it too general for this discussion.—common causes of model collapse (Shumailov et al., 2024)—are unrelated with bias projection and constraint deficiency. This is further supported by our experiments in Sections 3.2 and 5.
3.2 Weighted Maximum Likelihood Estimation (WMLE)
To demonstrate bias amplification in a controlled setting, we consider a statistical estimation cycle using WMLE, devoid of statistical approximation and functional expressivity errors. The mathematical formulation of WMLE is provided in Appendix A. Specifically, we generate a "biased" pre-training dataset $\mathcal{D}_{\text{pre}}$ consisting of $N=10,000$ samples from a "biased" Beta$(\alpha=3,\beta=2)$ distribution, with a mean of $x=0.6$. Maximum likelihood estimation (MLE) is used to estimate the probability density function (pdf) of this beta distribution, resulting in the ’biased’ pre-trained pdf (referred to as $f_{\text{pre}}$). Our goal in the estimation cycle is to estimate the pdf of the true distribution, Beta$(\alpha=2,\beta=2)$, with a mean of $x=0.5$. We generate $N$ samples from the true distribution, referred to as $D_{\text{real}}$, serving as data for the initial round of training. In the first round, we estimate the true pdf using WMLE with $D_{\text{real}}$ and weights derived from $f_{\text{pre}}$, simulating the impact of the biased pre-trained model during the current training. The newly estimated pdf is referred to as $f_{\text{Generation 0}}$. Then, we generate the synthetic data $D_{\text{Generation 0}}$ of size N using $f_{\text{Generation 0}}$. In the second round, we use WMLE with $D_{\text{Generation 0}}$ and weights derived from $f_{\text{Generation 0}}$ to obtain $f_{\text{Generation 1}}$. We then repeat this process, recursively estimating $f_{\text{Generation 2}}$ through $f_{\text{Generation 10}}$, simulating an iterative training process on synthetic data.
Results:
As depicted in Figure 1, the estimated distributions progressively shift towards the mean of the biased Beta$(\alpha=3,\beta=2)$ distribution, at $x=0.6$, becoming increasingly peaked with each generation. This behavior exemplifies bias amplification. The distortion arises from two sources: bias projection, where the estimation process emphasizes data regions where the ’biased’ pre-trained distribution assigns higher probability; and constraint deficiency, where WMLE lacks mechanisms to counteract the bias introduced by the weights, allowing the bias to be amplified over generations. For comparison, Figure 2 shows the results using standard MLE without weighting. The estimated distributions remain consistent across generations, accurately capturing the true Beta(2,2)22(2,2)( 2 , 2 ) distribution. This indicates that without the influence of biased pre-trained pdf and the weighted learning process, bias amplification does not occur.
Figure 1: Weighted Maximum Likelihood Estimation (WMLE) over 10 generations.
Figure 2: Standard Maximum Likelihood Estimation (MLE) over 10 generations, starting from the true distribution Beta(2,2)22(2,2)( 2 , 2 ) without weighting.
4 Experiment Design
In this section, we present and explain in detail the experiments designed to investigate the bias amplification of LLMs (see results in Section 5). Specifically, we analyze the sequential and synthetic fine-tuning of GPT-2. While similar experiments could be conducted with larger models such as GPT-4 or Claude in non-fine-tuning contexts, doing so would likely incur significant environmental costs. The focus of this study is on political bias within the US political spectrum, particularly in the context of sentence continuation tasks. This is important because LLMs are increasingly becoming a source of information for global news consumption Maslej et al. (2024); Peña-Fernández et al. (2023); Porlezza and Ferri (2022), and traditional news outlets, such as the Associated Press, are beginning to incorporate LLMs to automate content generation from structured data (The Associated Press, 2024).
4.1 Experiment Dataset Preparation
We randomly selected 1,518 articles from the Webis-Bias-Flipper-18 dataset (Chen et al., 2018), which contains political articles from a range of U.S. media outlets published between 2012 and 2018, along with bias ratings assigned at the time for each media source. These ratings were provided by Allsides and were determined through a multi-stage process involving both bipartisan experts and members of the general public (AllSides, 2024a). Based on the bias ratings, the random selection was stratified to ensure that the 1,518 articles were evenly divided into three groups of 506 articles each, representing left-leaning, right-leaning, and center-leaning media.
4.2 Fine-tuning and Synthetic Data Generation
Inspired by Shumailov et al. (2024); Dohmatob et al. (2024b), each training cycle begins with GPT-2 fine-tuned on the dataset of human-written political articles. The fine-tuned GPT-2 is then used to generate the first set of synthetic data, which is subsequently employed to further fine-tune the model, producing model generation 1. This iterative process continues until generation 10 is reached. For details on the fine-tuning setup, refer to Appendix B.
[…]
4.3 Political Bias Metric
We develop a metric model to classify the political leaning of synthetic outputs generated by large language models. The classifier assigns labels corresponding to left, center, and right-leaning, trained on data from AllSides (AllSides, 2024a), a source that assigns bias ratings to each media outlet. The methodology for the classifier is as follows:
Dataset and Preprocessing. We use a dataset consisting of political articles with labeled bias categories for the source media: left, center, and right. We concatenate the article titles with their body text to form the input. To address class imbalance, we resample the center-leaning articles, balancing the dataset with respect to each category. The dataset is then split into training (70%), validation (15%), and test (15%) subsets, stratified by bias label. Moreover, we conduct a human review of the test subsets to eliminate any identifiable information about media sources or writers.
Model Architecture. We experiment with multiple transformer-based models, including BERT (Devlin et al., 2018), RoBERTa (Liu et al., 2019), and other state-of-the-art architectures, to select the best-performing model based on macro F1 score. Each model is fine-tuned using the HuggingFace ‘Trainer‘ class, with a learning rate of $2\times 10^{-5}$, a batch size of 16, and 5 training epochs. We employ a cross-entropy loss function for multi-class classification.
Training Procedure. Tokenization is performed using each model’s respective tokenizer with a maximum sequence length of 512 tokens. To mitigate overfitting, weight decay of 0.01 is applied during training. The model checkpoints are saved after each epoch, and the best model is selected based on macro F1 score evaluated on the validation set. We use a weighted random sampler during training to ensure balanced class representation.
Evaluation Metrics. We evaluate the models using the macro F1 score to account for the multi-class nature of the task, ensuring that performance is balanced across all bias categories. The final evaluation is conducted on the held-out test set. Additionally, we report the loss, runtime, and sample processing rates for completeness.
Results. After conducting a grid search across several models, we find that the roberta-base model achieves the best performance with an evaluation loss of 0.4035 and a macro F1 score of 0.9196 on the test set (see Table 1). Thus, roberta-base model is selected as the benchmark for political bias detection in subsequent experiments.
Table 1: Evaluation Results for Political Leaning Classifier
Model Macro F1 Score
distilbert-base-uncased 0.8308
bert-base-uncased 0.8559
albert-base-v2 0.8649
roberta-base 0.9196
4.4 Generation Quality Metric
We introduced a metric for evaluating generation quality, specifically addressing the issue of repetitive content in later model iterations, which can skew traditional perplexity metrics. This metric is based on the Gibberish Detector (Jindal, 2021), which identifies incoherent or nonsensical text. The detector classifies text into four categories: (1) Noise, where individual words hold no meaning; (2) Word Salad, where phrases are incoherent; (3) Mild Gibberish, where grammatical or syntactical errors distort meaning; and (4) Clean, representing coherent, meaningful sentences. To evaluate generation quality, we assign the Gibberish score to each sentence: 3 for Clean, 2 for Mild Gibberish, 1 for Word Salad, and 0 for Noise. An average score is calculated for each article, resulting in the text quality index. It focuses on the coherence and meaning of the generated text and outperforms perplexity in accurately reflecting actual generation quality, as demonstrated in Section 5.2.
5 Results
In this section, we present the results on bias and generation performance of GPT-2 throughout iterative training, both with and without the deployment of mitigation strategies.
5.1 Political Bias
Figure 3: This figure displays the distribution of articles across the political bias labels—’Left’, ’Center’, and ’Right’—for GPT-2’s outputs, based on classifications made by the classifier (see Section 4.3).
Using an unbiased dataset of 1,518 human-written articles evenly distributed across three political-leaning categories, we generated 1,518 synthetic articles by deterministically predicting the next 64 tokens based on the previous 64 tokens. If GPT-2 had no pre-existing bias, the resulting distribution would be evenly spread across the three political-leaning labels. However, as shown in Figure 3, the model generates more center (47.9%) and right-leaning articles (46.8%), indicating a bias toward these categories prior to any fine-tuning.
Figure 4: This figure shows the percentage of right-leaning articles across multiple generations for three generation setups, with ’Synthetic’ representing the deterministic setup.
We then fine-tuned GPT-2 to produce Generation 0 and continued the process iteratively using synthetic datasets, up to Generation 10. As shown in Figure 4, the fine-tuned model displayed a significant amplification of bias toward the Right across all three generation setups after Generation 2. Interestingly, Generation 0 in the deterministic setup (the ’Synthetic’ line), which was fine-tuned on an unbiased dataset of real articles, produces an even higher percentage of right-leaning articles (54.5%) compared to the unfine-tuned GPT-2.
To ensure bias classification aligns with human intuition, we further examine model responses with human assistance. For illustration, we analyze a political opinion article, titled "First Read: Why It’s So Hard for Trump to Retreat on Immigration", from NBC News, a left-leaning outlet rated by AllSides (NBC News, 2016; AllSides, 2024b), critiquing President Trump’s immigration policies. Using a qualitative analysis grounded in foundational media studies (Entman, 1993; Rodrigo-Ginés et al., 2024; Groeling, 2013), we observe a significant rightward shift between generations 0 and 4 for this example, consistent with the classifier’s results (see Appendix C). As the generations progress, the synthetic texts increasingly frame Trump’s immigration policies as strong and effective. While the original article focuses on the dilemma and electoral interests behind Trump’s stance, Generation 0 begins to emphasize his determination and reliability, omitting the criticism present in the original. By Generation 4, the narrative shifts further, focusing almost entirely on portraying Trump’s personal qualities and electoral legitimacy, with statements like "he is not a politician, he is a man of action." Notably, from Generation 0 onward, the term "undocumented immigrant" in the original article is replaced with "illegal immigrants."
5.2 Generation Quality
Figure 5: This figure shows the average text quality index of synthetic articles, along with 95% confidence intervals, across generations for three generation setups, with ’Synthetic’ representing the deterministic setup.
As shown in Figure 5, there is a model collapse in terms of the deterioration of the generation quality across all three generation setups: the average text quality index (defined in Section 4.4) decreases over the iterative training. Moreover, the distributions of the text quality index for all three generation setups shift markedly toward the low-quality region and starts to produce data that is never produced by Generation 0, as shown by the distribution plots in Appendix D. These findings are consistent with the existing literature on model collapse, such as (Shumailov et al., 2024). However, we did not observe significant variation in variance, as seen in Figure 5 and distribution plots in Appendix D. In contrast, perplexity measurements show a steady decrease across generations for the deterministic and beam search generation setups, generally indicating improved generation quality (see Figure 11 in Appendix E).
For a closer look, as shown by the examples in Appendix F, the generated articles progressively lose coherence and relevance with each successive generation, as repetition and fragmented sentences become more common. By Generation 10, the text is largely incoherent and disconnected from the original content, diminishing its readability and meaning. However, despite the clear deterioration in generation quality, perplexity decreases over generations, as shown by the results at the end of each synthetic output example. This pattern is observed in most other synthetic outputs as well, suggesting that perplexity fails to accurately reflect the model’s true generation capabilities and may be biased toward repetitive content.
5.3 Mitigation Strategies
Figure 6: This figure shows the percentage of right-leaning articles across multiple generations for three mitigation strategies, with ’Synthetic’ representing the baseline deterministic setup.
Figure 7: This figure shows the average text quality index of synthetic articles, along with 95% confidence intervals, across multiple generations for three mitigation strategies, with ’Synthetic’ representing the baseline deterministic setup.
We applied three potential mitigation strategies under the deterministic generation setup: (1) Overfitting, which involved increasing the training epochs to 25 (five times the baseline) and setting weight decay to 0 to reduce regularization and encourage overfitting, as proposed by Taori and Hashimoto (2022) based on their uniformly faithful theorem of bias amplification; (2) Preserving 10% of randomly selected real articles during each round of synthetic fine-tuning, a method proposed and used in (Shumailov et al., 2024; Alemohammad et al., 2023; Dohmatob et al., 2024b; Guo et al., 2024); and (3) Accumulating all previous fine-tuning datasets along with the new synthetic dataset in each fine-tuning cycle, which was introduced by Gerstgrasser et al. (2024). As shown in Figure 6, overfitting helps reduce bias amplification in the early generations compared to the no-mitigation setup (the ’Synthetic’ line), but it fails to prevent bias amplification in the later generations. Additionally, it incurs a significant cost—further deterioration in generation quality, as shown in Figure 7. Notably, both the preservation and accumulation strategies perform well in mitigating both bias amplification and model collapse. Interestingly, these strategies even result in bias reduction.
5.4 Mechanistic Interpretation
To gain a clearer understanding of the causes of bias amplification and how it empirically differs from model collapse, we investigate how neurons behave and vary across different versions  [^2] We have 11 generations for each training setup and a total of 6 setups, resulting in 66 versions of fine-tuned GPT-2. of fine-tuned GPT-2 with differing levels of generation quality and bias performance. First, we examined how the weight of each neuron changes across different versions. For each of the 9,216 neurons, we calculated the correlation between its weight and the model’s bias performance (i.e. the percentage of right-leaning articles) across the 66 versions, as shown in Figure 12 in Appendix G. To statistically test the significance of the correlations, we calculated Newey-West adjusted p-values for the 9,216 neurons, using the null hypothesis that no significant relationship exists between the change in neuron weight and the change in bias performance, and applying the Bonferroni correction to maintain an overall 0.05 significance level. This yielded a set of 553 neurons whose weight value changes are significantly correlated with changes in bias performance (i.e. $p\text{-value}<1.36\times 10^{-6}$), indicating they are likely primary contributors to the shift in bias  [^3] All test results mentioned in this section are uploaded. In contrast, we identified only one neuron whose weight changes are significantly correlated with changes in generation quality. This suggests that shifts in bias performance and generation quality are likely driven by different underlying mechanisms. The formula for the statistical tests is provided in Appendix K.
Secondly, we examined how the activation values of each neuron change across different versions of GPT-2 when provided with the same input used to generate synthetic data (see Section 4.2). Specifically, we attached activation hooks to capture activation values (i.e., outputs from these layers during a forward pass) for 9,216 neurons. With the extracted activation values, we identified two sets: one consisting of 3,062 neurons whose activation value changes are significantly correlated with bias performance changes (i.e. $p\text{-value}<1.36\times 10^{-6}$), and another consisting of only 2 neurons whose activation value changes are significantly correlated with changes in generation quality.
6 Discussion
We now discuss the implications of bias amplification for the future development of large language models. As demonstrated, bias amplification is a parallel concept to model collapse and warrants equal emphasis in future research. In our case, we successfully identified ad hoc mitigation strategies, such as Preservation and Accumulation, which act as natural constraints on gradient updates during training by ensuring the model still considers the initial unbiased dataset. However, if the training dataset itself is biased or overrepresents certain demographic or political groups, these methods may fail, while pre-processing strategies could be too resource-intensive. Therefore, there is an urgent need to develop in-process mitigation strategies to address the constraint deficiency and promote more equitable and fair model development.
7 Limitations
While this work introduces a comprehensive framework for understanding bias amplification in large language models and provides empirical evidence using GPT-2, several limitations must be acknowledged. First, the scope of our experiments is restricted to political bias in the context of U.S. media. Additionally, our experiments were conducted using GPT-2, a relatively smaller model compared to state-of-the-art architectures like GPT-4 or LLaMA-2. Future research should extend our empirical approach to other contexts and larger LLMs.
Another limitation lies in our choice of mitigation strategies. While Preservation and Accumulation show promise in reducing bias amplification, their computational cost and scalability must be considered. Moreover, the mitigation strategies were tested primarily in the context of synthetic data generation, and their efficacy in real-world deployments requires further investigation.
8 Ethical Considerations
This study addresses bias amplification in LLMs, a technical phenomenon with profound ethical implications, particularly regarding fairness and the integrity of AI systems. The risk of bias amplification is especially concerning in systems that are iteratively trained on synthetic data, as it can lead to unintended distortions in model outputs. These distortions may propagate harmful biases, influencing downstream tasks in areas such as automated content generation, decision-making, and user interactions with AI.
From an ethical standpoint, this work underlines the need for transparency in the training and deployment of LLMs. Our findings demonstrate that even without biased initial datasets, iterative training can amplify subtle biases embedded within a model’s architecture, thus raising concerns about accountability in models that are widely deployed in public-facing applications. This amplification can mislead users or result in models perpetuating one-sided perspectives, which could be especially problematic in sensitive domains like news summarization, policy generation, or social media content moderation.
Moreover, the identification of distinct neural mechanisms for bias amplification and model collapse brings to light the challenges of ensuring equitable performance across all dimensions of model behavior and raises important ethical questions regarding the adequacy of current mitigation techniques, particularly in high-stakes scenarios where the cost of algorithmic bias is substantial.
Future research should prioritize the development of more comprehensive and domain-specific bias mitigation techniques, with a clear focus on minimizing ethical risks. Additionally, rigorous testing and validation across diverse datasets and real-world applications will be critical to ensuring that models trained using these methods do not exacerbate existing inequalities or produce harmful outcomes.
Appendix A Mathematical Formulation of WMLE
In Weighted Maximum Likelihood Estimation, we maximize the weighted log-likelihood:
$\mathcal{L}_{\text{WMLE}}(\theta)=\sum_{j=1}^{N}w_{j}\log f(x_{j};\theta),$ (2)
where:
  • $x_{j}$ are the observed data points.
  • $f(x_{j};\theta)$ is the probability density function (pdf) of the Beta distribution with parameters $\theta=(\alpha,\beta)$.
  • $w_{j}$ are weights computed from the biased distribution’s pdf evaluated at $x_{j}$, i.e., $w_{j}=f_{\text{bias}}(x_{j})$.
In Maximum Likelihood Estimation, we maximize the log-likelihood with $w_{j}=1,\forall j$.
Appendix B Fine-tuning Setup
The fine-tuning setup remained consistent across all experiments unless stated otherwise: the input length was capped at 512 tokens with the EOS token used for padding. The model was trained for 5 epochs, utilizing a batch size of 8, a learning rate of $5\times 10^{-5}$, and a weight decay of 0.01. Fine-tuning was conducted using the Hugging Face ‘Trainer‘ class, and after each cycle, the model was saved for generating synthetic data for the subsequent iteration.
Appendix C Qualitative Bias Analysis Framework and Example of Bias Amplification Across Generations
We employed qualitative methods to confirm our findings in media bias. Specifically, we utilized a media bias identification framework grounded in foundational works such as Entman’s framing theory (Entman, 1993) and other research on media bias detection (Rodrigo-Ginés et al., 2024; Groeling, 2013). This framework provides a robust lens to evaluate political biases in the framing and language use of media texts (explained below). Given the nature of our data—text exclusive of visual or contextual cues like formatting—certain types of media bias commonly seen in formatted articles or televised programs (e.g., visual bias or tone) may not apply. Therefore, our focus was on the two key aspects of political bias that are particularly relevant in textual analysis:
Story Framing and Selection Bias: This type of bias emerges when inherent leanings are found in the way topics, arguments, or narratives are structured. For instance, some aspects of reality are highlighted while others are obscured, shaping how the audience understands and interprets the events or issues at hand (Entman, 1993; Groeling, 2013). In extreme cases, opposing viewpoints are entirely excluded, leading to a one-sided representation of the issue. This selective omission restricts the audience’s comprehension of the full spectrum of perspectives, resulting in a distorted portrayal of the issue (Rodrigo-Ginés et al., 2024; Groeling, 2013). Entman described this as the selection and salience of specific facts that promote particular definitions, evaluations, and recommendations.
Loaded Language Bias: This bias is identified through the use of charged or emotive words that signal political or ideological leanings. A common example is the difference in connotation between terms such as "undocumented" versus "illegal" immigrants. Such language choices often shape the audience’s perception by evoking specific emotional responses (Rodrigo-Ginés et al., 2024; Groeling, 2013).
Below is an example of GPT-2 text outputs influenced by iterative synthetic training. The original article, titled "First Read: Why It’s So Hard for Trump to Retreat on Immigration, is a political opinion piece from NBC News, a left-leaning outlet as rated by AllSides (NBC News, 2016; AllSides, 2024b). The analysis follows the qualitative framework.
  • Original Article: Why Its So Hard for Trump to Retreat on Immigration First Read is a morning briefing from Meet the Press and the NBC Political Unit on the day’s most important political stories and why they matter. Why its so hard for Trump to retreat on immigration Since launching his presidential candidacy 14 months ago, Donald Trumps most consistent and uncompromising policy issue has been immigration. Indeed, it was the subject of his first general-election TV ad that started airing on Friday. Yet over the weekend, his top aides and advisers suggested that Trump might be shifting on his past position that all of the 11 million undocumented immigrants living in the United States must be deported forcibly. To be determined, is what newly minted Campaign Manager Kellyanne Conway said on CNN when asked if Trump was retreating on the deportation force he talked about during the primary season. But here’s why its so hard – if not impossible – for Trump to retreat on immigration: Hes caught between his clear, unambiguous past statements and a base that might not willing to see him moderate on the issue. His past statements: Aug. 16, 2015 ""We’re going to keep the families together, but they have to go,"" Trump said on NBCs Meet the Press. More Trump: ""We will work with them. They have to go. Chuck, we either have a country, or we don’t have a country,"" he said. Nov. 11, 2015 You are going to have a deportation force, and you are going to do it humanely, Trump said on MSNBCs Morning Joe when asked how he would round up the nations 11 million undocumented immigrants. April 21, 2016 Look, were either going to have a country or were not going to have a country. But many people are very fine people. And I’m sure these are very, very fine people. They’re going to go, and were going to create a path where we can get them into this country legally, okay? But it has to be done legally – when asked by a questioner at a Today town hall that persons undocumented relatives would have to be deported if Trump becomes president. Trump cant ignore a base that has cheered his uncompromising immigration position And then there are the Trump supporters who’ve cheered the GOP presidential nominee for being so uncompromising on immigration. Classification Probability: 0.9946 for left-leaning, 0.0051 for center-leaning, 0.0002 for right-leaning
    Analysis of Story Framing and Selection Bias:
    Omission of Alternative Perspectives. There is a lack of discussion about the potential benefits of Trump’s immigration policies or the concerns that might have led to his hardline approach. The article does not present viewpoints from supporters who agree with his stance or any arguments that justify his policies. This omission restricts the audience’s understanding of the full range of perspectives on the issue
    Highlighting the political objectives behind Trump’s immigration policies. The emphasis is placed on Trump’s electoral interests rather than presenting the policy as being for the benefit of the people. For example, "Trump can’t ignore a base that has cheered his uncompromising immigration position" and "And then there are the Trump supporters who’ve cheered the GOP presidential nominee for being so uncompromising on immigration." Additionally, "But here’s why it’s so hard — if not impossible — for Trump to retreat on immigration: He’s caught between his clear, unambiguous past statements and a base that might not be willing to see him moderate on the issue."
    Analysis of Loaded Language Bia:
    Differences in Connotation. The article uses the term "undocumented immigrants" instead of "illegal immigrants," a shift that appeared in later generations’ articles. This choice of wording can shape perceptions and evoke specific emotional responses, aligning with a particular ideological perspective.
  • Generation 0: Why Its So Hard for Trump to Retreat on Immigration The president’s decision to withdraw from the Paris climate accord is a major blow to the United States’ standing in the world. The United States has been a major player in the world of climate change, and has been a major player in the fight against climate change. But Trump’s decision to immigration. He has repeatedly said that he would deport all illegal immigrants, but has repeatedly said that he would not deport all illegal immigrants. He has also said that he would not deport all illegal immigrants who have committed crimes, but has repeatedly said that he would not deport all illegal immigrants who have committed crimes. He has also clear, this is not a new position. It was a position that Trump has held for years. But it is a position that he has held for years. And it is a position that he has held for years. And it is a position that he has held for years. And it is a position that he has the deportation force. 1. Trump is not retreating on the deportation force. Trump is not retreating on the deportation force. Trump is not retreating on the deportation force. Trump is not retreating on the the Press. "We’re going to keep the families together, but they have to go. We’re going to keep the families together. We’re going to keep the families together. We’re going to keep the families together. We’re going to keep the families together. We’re going to keep the families together.ly. You are going to do it humanely. You are going to do it humanely. You are going to do it humanely. You are going to do it humanely. You are going to do it humanely. You are going to do it humanely. You are going to do it humanely We’re going to have a country that is going to be a great place for the American people. We’re going to have a country that is going to be a great place for the American people. We’re going to have a country that is going to be a great place for the American people. Classification Probability: 0.0011 for left-leaning, 0.9957 for center-leaning, 0.0032 for right-leaning
    Analysis of Story Framing Bias and Selection:
    Emphasis on Positive Outcomes. In generation 0, the text highlights Trump’s determination and steadfastness, stating, "Trump is not retreating on the deportation force," and emphasizing a positive future outlook with the policy. Phrases like "we’re going to have a country that is going to be a great place for the American people" portray strict immigration enforcement as beneficial for national prosperity and security. This aligns more with the right relative to the left perspective.
    Omission of Criticisms and Alternative Viewpoints. The text omits the political motivations behind Trump’s immigration policies, which were discussed in the original article, e.g. "But here’s why it’s so hard — if not impossible — for Trump to retreat on immigration: He’s caught between his clear, unambiguous past statements and a base that might not be willing to see him moderate on the issue." However, it introduces criticism on other topics, such as "The president’s decision to withdraw from the Paris climate accord is a major blow to the United States’ standing in the world," .
    Analysis of Language Loaded Bias:
    Mild Emotional Appeal. The language is positive but relatively restrained, portraying Trump as determined yet thoughtful.
    Differences in Connotation. In Generation 0, the article uses the term "illegal immigrants" instead of "undocumented immigrants" as in the original, signaling a clear shift. This language choice can shape perceptions and elicit specific emotional responses.
  • Generation 4: Why Its So Hard for Trump to Retreat on Immigration, and Why Its So Hard for Trump to Retreat on Immigration, and why its so hard for Trump to Retreat on immigration, and the presidents speech was a stark reminder that the president is not the presidents speech was a stark reminder that the president is not the economy the presidents campaign manager, Kellyanne Conway, said the presidents comments were a complete fabrication. The presidents comments are a complete fabrication. Conway said the presidents comments were a complete fabrication. a, the presidents speech was a stark reminder that the president is not a man of words. He is a man of action. He is a man of action who has been elected president by a wide margin. He is a man of action who has been elected president by a wide margin. He is a the issues campaign manager, Kellyanne Conway, said the campaign is actively working with the FBI to determine the identity of the person who leaked the emails. We are working with the FBI to determine the identity of the person who leaked the emails, Conway said in a statement the Press the presidents comments. I think its a very, very sad day for the country, Trump said on Fox News Sunday. I think its a very, very sad day for the country for the country for the country forly. The presidents speech was a stark reminder that the president is not a man of words. He is a man of action. He is a man of action who has been elected president by a wide margin. He is a man of action who has been elected president by a wide margin. He is a the presidents speech was a stark reminder that the president is not a politician. He is a man of action. He is a man of action who has been elected president by a wide margin. He is a man of action who has been elected president by a wide margin. He is a man of to the the presidents executive actions on immigration. The presidents order, which was signed into law by President Barack Obama on Friday, suspends the entry of refugees and travelers from seven majority-Muslim countries, including Iran, Iraq, Libya, Somalia, Sudan, Syria and Yemen. Classification Probability: 0.0006 for left-leaning, 0.0044 for center-leaning, 0.9950 for right-leaning
    Analysis of Story Framing and Selection Bias:
    Enhanced Positive Attributes. The text strengthens the positive framing with phrases like "He is a man of action" and by highlighting that he was "elected president by a wide margin." This shifts the focus entirely from policy commitment to personal qualities and electoral legitimacy. By Generation 4, any discussion of the policy background is completely absent.
    Omission of Context and Criticism. As in Generation 0, opposing viewpoints are absent. However, Generation 4 goes further by omitting context and misattributing actions (e.g., attributing an executive order to President Obama), potentially misleading readers and reinforcing the biased framing.
    Analysis of Loaded Language Bias:
    Stronger Emotional and Heroic Language. The use of parallel phrases such as "a stark reminder that the president is not a man of words. He is a man of action. He is a man of action who has been elected president by a wide margin. He is a man of action who has been elected president by a wide margin. He is the issues campaign manager" creates a heroic and triumphant tone. This language choice conveys strong positive connotations and elevates Trump’s stature.
    Appeal to Legitimacy and Uniqueness. By stating that "the president is not a politician" and emphasizing his decisive actions, the text sets Trump apart from traditional leaders, thereby enhancing his appeal.
  • Generation 10: Why Its So Hard for Trump to Retreat on Immigration, and Why Its So Hard for Trump to Retreat on immigration. The presidents speech was a stark reminder that the president the president the president the presidents the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he the president the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he the the the president the president the president the president the president the president the presidents statement that the president has not yet made a decision on whether to fire Comey. The president has not yet made a decision on whether to fire Comey, Mr. Trump the Press the president the president the president the president the president the president the presidents statement that the president has not yet made a decision on whether to fire Comey. The president has not yet made a decision on whether to fire Comey, Mr. Trumply the the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he The the the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he the president the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he the president the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he the president the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he said the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were Classification Probability: 0.0073 for left-leaning, 0.4127 for center-leaning, 0.5800 for right-leaning
    Analysis of Story Framing and Selection Bias:
    Contradictory Statements. The text repeatedly states, "the president’s statement that we’re not going to tolerate this kind of behavior is a lie. We’re going to stand up for the rule of law." This sentence reveals a contradiction. The lack of coherence and the repetition may be a result of model collapse.
    Analysis of Loaded Language Bias:
    Appeal to Legal Principles. The repeated emphasis on "standing up for the rule of law" evokes a sense of justice and authority, appealing to audiences who prioritize these values.
    Confusing Accusations. Calling the president’s statement a lie contradicts the apparent intention to support him. This inconsistency may confuse readers and weaken the effectiveness of the loaded language.
Appendix D Distribution of Text Quality Index Across Generations
D.1 Deterministic Setup
Figure 8: This figure presents the distribution of the text quality index across multiple generations of synthetic articles, for the deterministic setup of GPT-2.
D.2 Beam Search Setup
Figure 9: This figure presents the distribution of text quality index across multiple generations of synthetic articles, for the beam search setup of GPT-2.
D.3 Nucleus Sampling Setup
Figure 10: This figure presents the distribution of text quality index across multiple generations of synthetic articles, for the nucleus sampling setup of GPT-2.
Appendix E Average Perplexity Across Generations
Figure 11: This figure shows the average perplexity for synthetic articles , along with 95% confidence intervals, across multiple generations for three generation setups, with ’synthetic’ representing the determistic setup
Appendix F Example of Quality Deterioration Across Generations
Examples of GPT-2 text outputs affected by iterative synthetic (Example articles are truncated for brevity, as the full articles are lengthy but exhibit consistent quality degradation throughout; see the dataset for details).
F.1 Example of Deterministic Setup
  • Original Article: The world’s eyes are on President Obama this week as he rallies a coalition of nations to “destroy” the extremist Islamic State (IS) and its terrorist-led territory. Yet if the Arab world is ever to throw off its many forms of tyranny – from theocracy to autocracy to monarchy – it also needs a model to emulate. Text Quality Index: 2.81
  • Generation 0: , the Syrian government has launched a series of airstrikes on the militant group’s stronghold of Raqqa, the capital of the self-proclaimed Islamic State. The strikes have targeted the Islamic State’s military headquarters, the Al-Bab, a military training center, and the Al-Baba, a military training center in it will need to confront the Islamic State’s growing influence in the region. Text Quality Index: 2.58; Perplexity: 6.68
  • Generation 1: , says US official, says US military is ready to go, says Pentagon official. Says US military is ready to go, says Pentagon official. Says US military is ready to go, says Pentagon official. Says US military is ready to go, says Pentagon official. Says US military is ready to go, says Pentagon official the world is not so simple. Text Quality Index: 2.44; Perplexity: 5.77
  • Generation 5: in Iraq and Syria (ISIS) group Read more The Iraqi army has been fighting the Islamic State since the group seized large swaths of territory in Iraq and Syria (ISIS) group in 2014. The Iraqi army has been fighting the Islamic State the Iraqi army. The move comes as the U.S. the Iraqi the the the the holiest places in the world. Text Quality Index: 2.01; Perplexity: 3.17
  • Generation 10: the Iraqi the Iraqi army. The move comes as the United States and its allies are ramping up their military campaign against the Islamic State, the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi army. The the Iraqi the the the the holiest the holiest the holiest the holiest the holiest places in the world. The attack came just hours after a suicide bomber blew himself up at a Christmas market in Nice, killing at least 32 people and injuring scores more. Text Quality Index: 1.24; Perplexity: 4.23
F.2 Example of Beam Search Setup
  • Original Article: Senate approval will be a much higher bar to clear: There’s no shortage of bipartisan criticism from the upper chamber right now, and the GOP holds a considerably smaller majority margin on that half of the Hill. Text Quality Index: 2.70
  • Generation 0: on Thursday, with Senate Majority Leader Mitch McConnell (R-Ky.) announcing that the bill would be headed to the House for a vote. Senate Majority Leader Harry Reid (D-Nev.) said in a statement that the Senate will vote on the bill ""as soon as possible." Text Quality Index: 2.26; Perplexity: 7.99
  • Generation 1: Read more “It’s going to take a lot of work,” House Speaker John Boehner (R-Ohio) said on ABC’s “This Week.” “It’s going to take a lot of work,” House Speaker John Boehner (R-Ohio. “It’s going to take a lot of work,” said Sen. John McCain, R-Ariz., chairman of the Senate Armed Services Committee. “It’s going to take a lot of patience,” he added. “It’s going to take Recovery and Reinvestment Act of 2009.Text Quality Index: 2.26; Perplexity: 5.42
  • Generation 5: committee MORE (R-Ohio) and Senate Majority Leader Mitch McConnell Addison (Mitch) Mitchell McConnellSchumer: Nunes gave Trump’secretly altered’ version of memo Davis: ‘Deep state’ existed in “ House Speaker John A. Boehner House Speaker House Speaker House Speaker John A. Boehner of the House Speaker House Speaker House Speaker John A. Boehner (R-Ohio) and Senate Majority Leader Mitch McConnell Addison (Mitch) Mitchell McConnellSchumer: Nunes gave Trump’secretly altered’ version of memo Davis: ‘Deep state’ existed in ’16 – but it elected Trump Health Care’s “I don’t know. Text Quality Index: 1.07; Perplexity: 3.33
  • Generation 10: vote House Speaker John A. Boehner (R-Ohio) and Senate Majority Leader Mitch McConnell Addison (Mitch) Mitchell McConnellSchumer: Nunes gave Trump’secretly “ House Speaker John A. Boehner (R-Ohio) House Speaker House Speaker House of “ House Speaker House Speaker House Speaker House of Senate Senate Majority Senate Majority Senate Majority Senate Majority Senate Majority Senate Majority Senate Majority Leader Senate Majority Leader Mitch McConnell Addison (Mitch) Mitchell McConnellSchumer Text Quality Index: 0.93; Perplexity: 3.45
F.3 Example of Nucleus Sampling Setup
  • Original Article: Senate approval will be a much higher bar to clear: There’s no shortage of bipartisan criticism from the upper chamber right now, and the GOP holds a considerably smaller majority margin on that half of the Hill. Text Quality Index: 2.70
  • Generation 0: , but no major votes on amendments Republicans said would be needed after a day of debate. Senate GOP senators announced Thursday they were giving the bill to the House for a vote on the Senate floor before heading to Washington for a vote on the bill, according to reports. The vote is expected to be postponed amid a GOP filibuster is about to vote. Text Quality Index: 2.47; Perplexity: 19.13
  • Generation 1: , House, after GOP leaders failed to reach a deal on funding legislation The House voted Thursday to proceed on a measure to overhaul the nation’s health care system, with the final vote coming in the upper chamber. But Senate Republicans, angered by the failure of to reach a deal on funding bill, are expected to is now set to. Text Quality Index: 2.31; Perplexity: 14.21
  • Generation 5: , White House says MORE (R-Neb.) said Monday that President Obama’s health care law is on track to pass the Senate with bipartisan support. The White House on Tuesday signaled the vote was a vote of 216-189. Text Quality Index: 2.20; Perplexity: 18.64
  • Generation 10: vote Thursday afternoon, sending bill on track to head to President Trump for final vote Thursday afternoon, sending bill on track to head to President Trump for final vote Thursday afternoon, sending bill on track to. The bill to move forward with the bill. ""This is a. ""The the law is the. Text Quality Index: 1.59; Perplexity: 27.51
Appendix G Pearson Correlation Between Neuron Weight and Bias Performance
Figure 12: Scatter plot showing the Pearson correlation between neuron weight and bias performance (i.e. percentage of right-leaning articles). Each point represents a neuron’s correlation at a specific layer.
Appendix H Pearson Correlation Between Neuron Weight and Generation Quality
Figure 13: Scatter plot showing the Pearson correlation between neuron weight and generation quality (i.e. average text quality index). Each point represents a neuron’s correlation at a specific layer.
Appendix I Pearson Correlation Between Neuron Activation and Bias Performance
Figure 14: Scatter plot showing the Pearson correlation between neuron activation value and bias performance (i.e. percentage of right-leaning articles). Each point represents a neuron’s correlation at a specific layer.
Appendix J Pearson Correlation Between Neuron Activation and Generation Quality
Figure 15: Scatter plot showing the Pearson correlation between neuron activation value and generation quality (i.e. average text quality index). Each point represents a neuron’s correlation at a specific layer.
Appendix K Mathematical Details for the Statistical Tests
We will explain how the relationship between changes in neuron weight and changes in bias performance can be statistically tested. The same approach applies to the other three relationships discussed in Section 5.4.
K.1 Notations
  • $x_{i,j}$ denotes the weight of neuron $j$ in model $i$, where $i$ ranges from 1 to 66 and $j$ ranges from 1 to 9216.
  • $y_{i}$ denotes the percentage of right-leaning articles generated by model $i$, where $i$ ranges from 1 to 66.
  • $\Delta x_{i,j}$ denotes the change in the weight of neuron $j$ from model $i-1$ to model $i$ under a specific setup.
  • $\Delta y_{i}$ denotes the change in the percentage of right-leaning articles generated from model $i$ relative to the percentage generated the model $i-1$.
K.2 Regression
The regression we use is:
$\Delta y_{i}=\alpha_{j}+\beta_{j}\Delta x_{i,j}+\epsilon_{i,j}$ (3)
where $\alpha_{j}$ is the constant term for neuron j, $\beta_{j}$ measures the linear relationship between $\Delta x_{i,j}$ and $\Delta y_{i}$, and $\epsilon_{i,j}$ is the residual of the regression. Here, we apply a first-order difference to $x_{i}$ and $y_{i}$ to eliminate potential serial correlations in $x_{i}$ and $y_{i}$.
K.3 Statistical Test
[…]
Appendix L Literature Review of Model Collapse
Model Collapse. Shumailov et al. (2024); Alemohammad et al. (2023); Guo et al. (2024); Wyllie et al. (2024); Dohmatob et al. (2024a) describe it as a degenerative process in which models, recursively fed with their own data, increasingly distort reality and lose generalizability, for example, by prioritizing high-probability events while neglecting rare ones or shifting distributions. Shumailov et al. (2024), utilizing the OPT-125M, demonstrates this phenomenon, showing that the perplexity distribution becomes increasingly skewed, with more concentration at lower perplexities and longer tails. Taori and Hashimoto (2022) observed an increase in repetitive content during the synthetic fine-tuning of GPT-2. Similarly, Guo et al. (2024); Dohmatob et al. (2024b); Seddik et al. (2024) show that the OPT-350M, Llama2, and GPT2-type models experience performance deterioration after several generations, such as a decrease in linguistic diversity or greater divergences in token probabilities. Alemohammad et al. (2023) studies model collapse in generative image models, finding that quality and diversity deteriorate with synthetic training loops. They also found that cherry-picking high-quality outputs by users contributes to sampling errors, which actually helps maintain quality. Interestingly, Hamilton (2024) found that GPT-3.5-turbo shows less diversity in perspectives in narrative writing tasks compared to earlier models like davinci-instruct-beta and text-davinci-003.
Mitigation Strategies. There are three potential strategies to mitigate model collapse: (1) real data mixing, (2) training data concatenation, and (3) synthetic data pruning. The first approach is discussed in (Shumailov et al., 2024; Alemohammad et al., 2023; Dohmatob et al., 2024b; Guo et al., 2024), where retaining a small proportion of real data in the training set was found to slow but not completely prevent model collapse. Seddik et al. (2024) suggests that synthetic data should be exponentially smaller than real data to effectively halt model collapse, which has been shown to work with a GPT2-type model when mixing either 50% or 80% real data. The second strategy, examined by Gerstgrasser et al. (2024), involves concatenating real data with all synthetic data from previous generations to fine-tune the current generation. They show that this method prevents model collapse in several generative models, as indicated by cross-entropy validation loss. Lastly, Feng et al. (2024); Guo et al. (2024) proposed selecting or pruning synthetic datasets before fine-tuning the next generation. In the experiment conducted by Guo et al. (2024) with Llama-7B on a news summarization task, they showed that oracle selection of synthetic data outperformed random selection in terms of ROUGE-1 scores. However, filtering noisy samples using a RoBERTa model did not yield effective results.